Max and Rebecca are in cars 120 miles apart. Both drivers are going at a constant speed of 30/mph and 40/mph respectively. How far apart will the 2 cars be in exactly 1 hour before they meet?
Answer by stanbon(75887)

Max and Rebecca are in cars 120 miles apart. Both drivers are going at a constant speed of 30/mph and 40/mph respectively. How far apart will the 2 cars be in exactly 1 hour before they meet?
-------
Determine when they will meet.
Max DATA:
rate = 30 mph ; time = t hrs ; distance = 30t miles
-----
Reb DATA:
rate = 40 mph ; time = t hrs ; distance = 40t miles
-----
Equation:
30t + 40t = 120 miles
70t = 120 miles
time = 12/7 hr.
------
One hour before = 5/7 hr
---
Max distance after 5/7 hr = (5/7)30 = 150/7 = 21.4 miles
Reb distance after 5/7 hr = (5/7)40 = 200/7 = 28.57 miles
------
Ans: Distance separating the cars:: 120-(21.4+28.57) = 70 miles
------------
Cheers,
Stan H.
-----------